find the slope of the line that passes throught the ordered pairs:
(a,b); (b,a)
(a,b); (-b,-a)
simplifiy and express your answer as a real number.

Answer by wgomero(28)

slope= (b - a)/(a - b)= - 1, les asume that b=4 and a=7, then (4 - 7)/(7 - 4) =
- 3/3= -1


slope = (b -(-a))/ (a -(-b)) (b + a )/( a + b) = 1
les asume that b=4 and a=7,
(4 + 7)/(7 + 4) = 1